A convex pentagon has three right angles. The other two angles are congruent to each other. What is the measure, in degrees, of one of these other angles?
Answer by Theo(13342)

sum of the interior angles of a convex pentagon = 540
3 * 90 = 270
540 - 270 = 270
270 / 2 = 135
-----
The other 2 angles are 135 degrees each.
3 * 90 + 2 * 135 = 270 + 270 = 540
-----